OpenStudy (ars22):

how do you find the zeros of a polynomial in y = ax^2 + bx + c ?

jhonyy9 (jhonyy9):

x_1,_2=(-b+/- sqrt(b2-4ac))/2a

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

factor things out so that youe left with something like (hx + d)(gx + e). set each thing in the ( ) = to 0, like (hx + d) = 0, then solve for x. x^2 + 3x + 2 = 0 (x + 1)(x + 2) = 0 (x + 1) = 0, (x + 2) = 0 x = -1, and -2